# Model-driven Stochastic Trace Clustering

### v1

@@ Abstract
Process discovery algorithms automatically extract process models from event logs, but high variability often results in complex and hard-to-understand models. To mitigate this issue, trace clustering techniques group process executions into clusters, each represented by a simpler and more understandable process model. Model-driven trace clustering improves on this by assigning traces to clusters based on their conformity to cluster-specific process models. However, most existing clustering techniques rely on either no process model discovery, or non-stochastic models, neglecting the frequency or probability of activities and transitions, thereby limiting their capability to capture real-world execution dynamics. We propose a novel model-driven trace clustering method that optimizes stochastic process models within each cluster. Our approach uses entropic relevance, a stochastic conformance metric based on directly-follows probabilities, to guide trace assignment. This allows clustering decisions to consider both structural alignment with a cluster’s process model and the likelihood that a trace originates from a given stochastic process model. The method is computationally efficient, scales linearly with input size, and improves model interpretability by producing clusters with clearer control-flow patterns. Extensive experiments on public real-life datasets show that our method outperforms existing alternatives in representing process behavior and reveals how clustering performance rankings can shift when stochasticity is considered.
@@ 1 Introduction
Cluster analysis is a fundamental technique in data mining and machine learning, aimed at grouping entities such that those within the same cluster are more similar to each other than to those in different clusters. In process mining, trace clustering applies this principle to event logs by grouping process executions (cases) into smaller, more coherent subsets. This is particularly valuable for complex processes, as it enables the discovery of simpler, more interpretable process models for each cluster. Such models enhance the readability of discovered process models and can also improve the accuracy of conformance checking [1, 2, 3]. However, many existing trace clustering methods are unsupervised and optimize generic similarity objectives that are not explicitly aligned with improving process model quality or interpretability [1]. Model-driven trace clustering addresses this by iteratively discovering a process model for each cluster and assigning traces based on how well they conform to these models [1, 4].
@@
Despite these developments, existing model-driven clustering methods have yet to incorporate the stochastic characteristics of event logs. To our knowledge, no approach currently integrates stochasticity into the clustering process to produce stochastic process models per cluster. To address this gap, we introduce a novel model-driven trace clustering method that incorporates stochastic conformance measuring into clustering decisions. Our method leverages entropic relevance (ER) [5], a stochastic conformance measure that quantifies the compression cost of traces in an event log based on directly follows probabilities. By minimizing ER within each cluster, the algorithm optimizes the representativeness of that cluster’s Directly-Follows Graph (DFG)—one of the most widely used model types in (commercial) process mining tools [6].
Our approach is computationally efficient, scaling linearly with the size of the event log, and supports integration with edge frequency filtering to mitigate the “spaghetti model” problem, potentially still present in downstream clusters. By promoting clusters with shared, high-probability directly-follows relations, it enhances model clarity and interpretability. Through extensive experiments on publicly available real-world datasets, we demonstrate that our method effectively captures control-flow behavior and yields interpretable, representative stochastic models. It also reveals that incorporating stochasticity can significantly alter the relative performance of clustering algorithms, underlining the importance of probabilistic considerations in trace clustering.
@@
The remainder of this paper is structured as follows: Section 2 introduces key terminology and discusses relevant background literature. Section 3 explains the workings of the entropic clustering algorithm. The experimental setup is detailed in Section 4, followed by the results presented in Section 5. Section 6 discusses these results along with potential limitations of the current approach. Finally, Section 7 provides a conclusion and outlines possible directions for future research. The code and the full results can be found online [^1] https://github.com/jaripeeperkorn/EntropicClustering.
2 Background
In this section, we provide an overview of process mining and its subdomain of trace clustering.
2.1 Event data and traces
The main data source used in process mining exists in the form of event logs, which record the executions of different activities within a business process. Each event typically includes a case ID that identifies the specific process instance it belongs to, an activity type that indicates what action was performed, and a timestamp or another attribute that defines the event’s position in the sequence. The ordered set of events for a case is referred to as a trace. Since the same sequence of activities may occur in multiple cases, an event log is formally a multiset of traces. In this work, we focus solely on control-flow information, meaning we represent traces as sequences of activity labels, disregarding other event attributes such as resources or timestamps. The set of all unique traces observed in an event log is referred to as the set of (control-flow) variants, with each variant’s frequency denoted by its multiplicity. A variant log is an abstraction of the full event log, capturing all distinct variants and their counts. To analyze behavior, traces are often compared to a process model such as a Petri net, which graphically depicts the logical structure and dynamics of a business process.
@@ 2.2 Process discovery & conformance checking
Process models can be manually designed or automatically discovered from event logs through process model discovery techniques, which aim to generate models that reflect and generalize the observed behavior. Well-known methods for automatically discovering Petri Nets include Heuristic Miner [7] and Inductive Miner [8], while other methods focus on other model types such as Fuzzy Miner [9] that can be used to create graph representations of the process at different levels of detail. Conformance checking techniques are used to compare an event log with a given process model, along different dimensions. Fitness measures how well the behavior found in the event log is described by the model, while precision measures how much of the allowed behavior is actually present in the event log. Typical fitness and precision metric include token-replay [10] and alignments [11, 12]. Generalization metrics aim to measure the model’s usefulness on unseen (but correct) process behavior. Simplicity reflects the readability, understandability, and complexity of the process, measured for example by the number of arcs [13]. Stochastic process models extend traditional process models by incorporating probabilities (or other quantitative aspects) of process executions. They provide insights into the likelihood of different behaviors. Stochastic conformance checking techniques incorporate these probabilities, e.g., by analyzing the statistical properties of execution patterns [14, 15, 16]. A widely used model type is the directly-follows graph (DFG), where nodes correspond to activity types and directed edges represent directly-follows (DF) relations observed in the log. DFGs are commonly used in commercial tools to visualize the process executions found in an event log [6]. By normalizing the edge frequencies with respect to the counts of their source activities, DFGs can also be interpreted as probabilistic (stochastic) models, with each DF relation assigned a transition probability. This property allows us to treat DFGs as stochastic process models in our work.
@@ 2.3 Trace clustering
Trace clustering refers to the task of grouping similar traces together based on control-flow patterns or other attributes. Clustering can help uncover subprocesses or behavioral variants, facilitating the discovery of simpler and more interpretable models for each group. A common line of work relies on vector-based methods, which transform traces into numerical feature vectors and apply standard clustering algorithms [2, 3, 17]. Context-aware approaches extend this by incorporating additional information about the execution context or structural dependencies [18]. A more process-centric alternative is provided by model-based clustering techniques. Ferreira et al. [4] proposed a method that learns a mixture of first-order Markov models using the Expectation-Maximization algorithm, but its computational cost limits scalability. ActiTraC [1] is another model-based approach that iteratively mines models using the Heuristic Miner algorithm [7] and evaluates the assignment of traces to clusters using fitness scores. Other works make use of alignments to perform trace clustering in different ways as well: using local alignments in combination with k-means clustering [19], by using model traces as centroids [20], or by explicitly using an available process model [21]. Despite existing advances, no method has yet integrated stochastic conformance information into the clustering process. In particular, model-driven clustering approaches have not leveraged the probabilistic behavior captured by stochastic process models. This gap motivates our contribution: a trace clustering technique that uses entropic relevance to evaluate and guide trace assignments.
@@ 3.1 Entropic Relevance
We use entropic relevance (or ER), a stochastic conformance checking measure that quantifies the information-theoretic compression cost of traces in an event log, given the structure and likelihoods of a stochastic process model [5]. A lower compression cost (i.e., a lower ER score) indicates a better fit between the model and the event log. For each trace, the ER contribution is computed using the probability $p$ assigned to that trace by the model, calculated as: $-\text{log}_{2}(p)$. The total ER score for a given log-model combination is obtained by summing these contributions over all traces in the log. For non-fitting traces, a so-called background model is used, giving a compression cost that is completely based on the log only (and is generally much higher). ER gives a score that measures quality along both the fitness and precision dimension. By punishing for a lack of fitness through the background model, or from a stochastic point of view, lower values for $p$ are assigned to highly unlikely cases. If the model allows too much behavior (with high probability), i.e., it is imprecise, the probabilities of the fitting cases will drop as well, leading to an increase in ER score.
In our approach, we use DFGs, which can be converted to stochastic process models. The discovery of a DFG can be done fast and efficiently, by just counting how often each activity appears and the frequency with which one activity directly follows another in the traces of the event log. For our purposes, the DFG is stored as two dictionaries: one containing the activity types (nodes) and their counts, and one the directly-follows relations (edges) and their counts. Given a DFG, the probability of a specific trace is computed as the product of the transition (edge) probabilities along the trace. Each transition probability is calculated by dividing the count of the corresponding edge by the count of its source node. In the original ER implementation, stochastic process models are converted into stochastic deterministic finite automata (SDFAs) to compute trace probabilities. To replicate this in the DFG-based approach, special Beginning-Of-Sequence (BOS) and End-Of-Sequence (EOS) activities are introduced to all traces and included in the DFG. Incorporating these artificial transitions effectively replicates the initialization and termination probabilities that would otherwise be handled in an SDFA. The traces with the added BOS and EOS activities are also the ones used to discover the DFG, i.e., the DFGs will also include those nodes. Since all traces now start with BOS and end with EOS, the transition probabilities from the BOS node to any other node can be seen as initialization probabilities, and those from each node to EOS as termination probabilities. Our (simplified) ER calculation is formalized in Algorithm 1. In order to be able to compare ER scores across different clusters (or event logs) with different amounts of traces, we also opted to use the average ER score over traces, rather than the sum over all traces.
@@
0: Variant log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$: $K$ distinct variants $t_{i}$, each with multiplicity $m_{i}$ directly-follows graph $G=(N,E,c_{N},c_{E})$: with $N=$ set of activity labels incl. BOS and $E=$ set of directed edges $c_{N}$ and $c_{E}$ the count of occurrences of the nodes and edges (the DFG). underflow threshold $\varepsilon$
@@
1:$T\leftarrow\sum_{i=1}^{K}m_{i}$Total number of trace occurrences}
2:$\mathrm{ER}_{\text{sum}}\leftarrow 0$Sum of per-trace information costs}
@@
13:    $\mathrm{ER}_{\text{sum}}\leftarrow\mathrm{ER}_{\text{sum}}-m_{i}\,\log_{2}(\hat{p})$$m_{i}$: multiplicity of $t_{i}$}
@@
15:$\overline{\mathrm{ER}}\leftarrow\mathrm{ER}_{\text{sum}}/T$Divide by total traces}
@@
Additionally, as mentioned above, the original ER formulation also handles non-fitting traces using a background model. However, in our method, ER scores are computed only for trace–DFG combinations in which the trace was used to discover the DFG. Thus, the model is guaranteed to fit the trace and its probability can be directly derived from the DFG without a background model. Finally, to avoid numerical underflow when computing probabilities for complex models (e.g., traces with many loops), the implementation constrains each trace probability to be at least $10^{-10}$ before taking the logarithm.
@@ 3.2 Entropic Clustering
Since ER computation scales linearly with input size [5], the clustering process (excluding initialization) is also linear in complexity with the exact runtime depending on trace lengths, the preselected number of clusters, and the number of unique variants. An alternative approach, which adds the variant to the cluster for which the total cluster ER score is minimal (after adding), rather than the specific variant’s ER, was also explored. However, initial results showed inferior performance, leading us to omit it from further experiments.
@@
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$, where $t_{i}$ is a trace variant and $m_{i}$ its multiplicity; number of clusters $k$
@@
5:    Build DFG $G_{j}$ from $s_{j}$
6:end for
7:for each variant $v\in\mathcal{V}\setminus\{s_{1},\ldots,s_{k}\}$, in decreasing order of multiplicity do
8:    for each cluster $C_{j}\in\mathcal{C}$ do
9:       Temporarily add $v$ to $C_{j}$ and update DFG $G_{j}$
10:       Compute average entropic relevance $\mathrm{ER}(v,G_{j})$
11:    end for
12:    Let $j^{*}=\arg\min_{j}\mathrm{ER}(v,G_{j})$
13:    Assign $v$ to cluster $C_{j^{*}}$ and update $G_{j^{*}}$ accordingly
14:end for
15:return  Final cluster assignment $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$
Figure 1: Illustrative example how the method selects the cluster.
Next to this, we also implemented an alternative variant of EC that iteratively splits the cluster with the highest ER score until the required number of clusters is reached (Algorithm 3). The motivation behind this approach is that by continuously refining the worst-performing cluster, we may achieve overall better results, similar to decision trees employing ID3 or CART.
Algorithm 3 Entropic Clustering Split (EC-Split)
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$, desired number of clusters $k$
@@
1: Initialize clustering $\mathcal{C}\leftarrow\text{EC}(L,2)$Run Entropic Clustering with 2 clusters}
@@
4:       Compute average entropic relevance $\overline{\mathrm{ER}}(C_{j})$
5:    end for
6:    Let $j^{*}=\arg\max_{j}\overline{\mathrm{ER}}(C_{j})$Identify worst-fitting cluster}
7:    Let $L_{j^{*}}$ be the variant log of cluster $C_{j^{*}}$
8:    Replace $C_{j^{*}}$ with two sub-clusters $\{C_{j^{*}1},C_{j^{*}2}\}\leftarrow\text{EC}(L_{j^{*}},2)$
9:    Update $\mathcal{C}\leftarrow(\mathcal{C}\setminus\{C_{j^{*}}\})\cup\{C_{j^{*}1},C_{j^{*}2}\}$
10:end while
11:return  Final cluster assignment $\mathcal{C}$

### v2

@@ Abstract
Process discovery algorithms automatically extract process models from event logs, but high variability often results in complex and hard-to-understand models. To mitigate this issue, trace clustering techniques group process executions into clusters, each represented by a simpler and more understandable process model. Model-driven trace clustering improves on this by assigning traces to clusters based on their conformity to cluster-specific process models. However, most existing clustering techniques rely on either no process model discovery, or non-stochastic models, neglecting the frequency or probability of activities and transitions, thereby limiting their capability to capture real-world execution dynamics. We propose a novel model-driven trace clustering method that optimizes stochastic process models within each cluster. Our approach uses entropic relevance, a stochastic conformance metric based on directly-follows probabilities, to guide trace assignment. This allows clustering decisions to consider both structural alignment with a cluster’s process model and the likelihood that a trace originates from a given stochastic process model. The method is computationally efficient, scales linearly with input size, and improves model interpretability by producing clusters with clearer control-flow patterns. Extensive experiments on public real-life datasets demonstrate that while our method yields superior stochastic coherence and graph simplicity, traditional fitness metrics reveal a trade-off, highlighting the specific utility of our approach for stochastic process analysis.
@@ 1 Introduction
Cluster analysis is a fundamental technique in data mining and machine learning, aimed at grouping entities such that those within the same cluster are more similar to each other than to those in different clusters. In process mining, trace clustering applies this principle to event logs by grouping process executions (cases) into smaller, more coherent subsets. This is particularly valuable for complex processes, as it enables the discovery of simpler, more interpretable process models for each cluster. Such models enhance the readability of discovered process models and can also improve the accuracy of conformance checking [DeWeerdt_2013, Greco2006, Song2009]. However, many existing trace clustering methods are unsupervised and optimize generic similarity objectives that are not explicitly aligned with improving process model quality or interpretability [DeWeerdt_2013]. Model-driven trace clustering addresses this by iteratively discovering a process model for each cluster and assigning traces based on how well they conform to these models [DeWeerdt_2013, Ferreira2007].
@@
Despite these developments, existing model-driven clustering methods have yet to incorporate the stochastic characteristics of event logs. To our knowledge, no approach currently integrates stochasticity into the clustering process to produce stochastic process models per cluster. To address this gap, we introduce a novel model-driven trace clustering method that incorporates stochastic conformance measuring into clustering decisions. Our method leverages entropic relevance (ER) [Alkhammash2022], a stochastic conformance measure that quantifies the compression cost of traces in an event log based on directly follows probabilities. By minimizing ER within each cluster, the algorithm optimizes the representativeness of that cluster’s Directly-Follows Graph (DFG), one of the most widely used model types in (commercial) process mining tools [Vanderaalst2019]. This focus makes our method particularly advantageous in scenarios where understanding the probability of process paths is as critical as the structural flow itself.
Our approach is computationally efficient, scaling linearly with the size of the event log, and supports integration with edge frequency filtering to mitigate the “spaghetti model” problem, potentially still present in downstream clusters. By promoting clusters with shared, high-probability directly-follows relations, it enhances model clarity and interpretability. Through extensive experiments on publicly available real-world datasets, we demonstrate that our method effectively captures control-flow behavior and yields interpretable, representative stochastic models. The results also reveal that incorporating stochasticity significantly alters the relative performance of clustering algorithms, suggesting that Entropic Clustering is the preferred choice for tasks requiring high stochastic precision, whereas traditional methods may still suffice for purely structural conformance.
@@
The remainder of this paper is structured as follows: Section 2 introduces key terminology and Section 3 discusses relevant background literature. Section 4 explains the workings of the entropic clustering algorithm. The experimental setup is detailed in Section 5, followed by the results presented in Section 6. Section 7 discusses these results along with potential limitations of the current approach. Finally, Section 8 provides a conclusion and outlines possible directions for future research. The code and the full results can be found online [^1] https://github.com/jaripeeperkorn/EntropicClustering.
2 Preliminaries
Event logs record executions of business process activities. Each event typically includes a case identifier (case ID) that identifies the process instance it belongs to, an activity label indicating which action was performed, and a timestamp or another attribute that defines the event’s position in the sequence. The sequence of events for a case is referred to as a trace. Let $\mathcal{A}$ be a finite set of activity labels. A trace $\sigma$ is then a finite sequence of such labels, that is, $\sigma\in\mathcal{A}^{*}$, and we denote its length by $|\sigma|$. An event log $L$ can be regarded as a multiset of traces, where each trace represents a process instance (case). Formally, $L$ is defined as a mapping $L:\mathcal{A}^{*}\to\mathbb{N}$, where for any $\sigma\in\mathcal{A}^{*}$, $L(\sigma)$ denotes the multiplicity (frequency) of $\sigma$ in the log. In this work, we focus solely on control-flow information and represent traces as sequences of activity labels, disregarding other event attributes such as resources or timestamps. The set of all unique traces in an event log is called the set of (control-flow) variants, with each variant’s frequency denoted by its multiplicity. A variant log is an abstraction of the full event log that captures all distinct variants and their counts.
To analyze behavior, traces are often compared with a process model that graphically depicts the logical structure and dynamics of a business process. Process models can be manually designed or automatically discovered from event logs through process model discovery techniques, which aim to generate models that reflect and generalize the observed behavior. Stochastic process models extend traditional process models by incorporating probabilities (or other quantitative aspects) of process executions, thereby providing insights into the likelihood of different behaviors. While various notations exist (e.g., Petri nets [Petri]), in this work, we mainly use the Directly-Follows Graph (DFG) as our modeling formalism. A DFG is a directed graph $G=(N,E)$, where $N\subseteq\mathcal{A}$ is the set of nodes (activities) and $E\subseteq N\times N$ is the set of edges representing directly-follows (DF) relations observed in the log. The directly-follows count (DF) between two activities $a$ and $b$, denoted $|a\to b|$, is the total number of times activity $a$ is immediately followed by $b$ in the cases. DFGs are commonly used in commercial tools to visualize the process executions found in an event log [Vanderaalst2019].DFGs can be turned into stochastic process models by annotating edges with transition probabilities. The transition probability is then defined as:
$P(b\mid a)=\frac{|a\to b|}{\sum_{x\in N}|a\to x|}$
This property allows us to treat DFGs as stochastic models, where the probability of a trace can be computed as the product of the transition probabilities of its sequence of activities. Trace clustering refers to the task of grouping similar traces based on control-flow patterns or other attributes. Clustering can help uncover subprocesses or behavioral variants, facilitating the discovery of simpler and more interpretable models for each group.
3 Related work
@@ 3.1 Process discovery & conformance checking
Well-known methods for automatically discovering Petri Nets include Heuristic Miner [Weijters2011] and Inductive Miner [Leemans2014], while other methods focus on other model types, such as Fuzzy Miner [Gunther2007] that can be used to create graph representations of the process at different levels of detail. Conformance checking techniques are used to compare an event log with a given process model, along different dimensions. Fitness measures how well the behavior found in the event log is described by the model, while precision measures how much of the allowed behavior is actually present in the event log. Typical fitness and precision metrics include token-replay [rozinat2008] and alignments [Vanderaalst2012, Adriansyah2015]. Generalization metrics aim to measure the model’s usefulness on unseen (but correct) process behavior. Simplicity reflects the readability, understandability, and complexity of the process, measured, for example, by the number of arcs [SanchezGonzalez2010]. Stochastic conformance checking techniques incorporate the probabilities found in stochastic process models, e.g., by analyzing the statistical properties of execution patterns [Leemans2019, Leemans2020, Leemans2023].
@@ 3.2 Trace clustering
A common line of trace clustering work relies on vector-based methods, which transform traces into numerical feature vectors and apply standard clustering algorithms [Greco2006, Song2009, DeKoninck2018]. Context-aware approaches extend this by incorporating additional information about the execution context or structural dependencies [Bose2010]. A more process-centric alternative is provided by model-based clustering techniques. Ferreira et al. [Ferreira2007] proposed a method that learns a mixture of first-order Markov models using the Expectation-Maximization algorithm, but its computational cost limits scalability. ActiTraC [DeWeerdt_2013] is another model-based approach that iteratively mines models using the Heuristic Miner algorithm [Weijters2011] and evaluates the assignment of traces to clusters using fitness scores. Other works make use of alignments to perform trace clustering in different ways as well: using local alignments in combination with k-means clustering [Evermann2016], by using model traces as centroids [Chatain2017], or by explicitly using an available process model [Boltenhagen2019]. Despite existing advances, no method has yet integrated stochastic conformance information into the clustering process. In particular, model-driven clustering approaches have not leveraged the probabilistic behavior captured by stochastic process models. This gap motivates our contribution: a trace clustering technique that uses entropic relevance to evaluate and guide trace assignments, offering a distinct alternative that prioritizes stochastic coherence over purely structural alignment.
@@ 4.1 Entropic Relevance
We use entropic relevance (or ER), a stochastic conformance checking measure that quantifies the information-theoretic compression cost of traces in an event log, given the behaviour of a stochastic process model as induced by its structure and transition probabilities [Alkhammash2022]. A lower compression cost (i.e., a lower ER score) indicates a better fit between the model and the event log. For each trace, the ER contribution is computed using the probability $p$ assigned to that trace by the model, calculated as: $\text{log}_{2}(p)$. The total ER score for a given log-model combination is obtained by summing these contributions over all traces in the log. For non-fitting traces, a so-called background model is used, giving a compression cost that is completely based on the log only (and is generally much higher). ER gives a score that measures quality along both the fitness and precision dimensions. By punishing for a lack of fitness through the background model, or from a stochastic point of view, lower values for $p$ are assigned to highly unlikely cases. If the model allows too much behavior (with high probability), i.e., it is imprecise, the probabilities of the fitting cases will drop as well, leading to an increase in ER score.
In our approach, we use DFGs, which can be converted to stochastic process models. The discovery of a DFG can be done fast and efficiently, by just counting how often each activity appears and the frequency with which one activity directly follows another in the traces of the event log. For our purposes, the DFG is stored as two dictionaries: one containing the activity types (nodes) and their counts, and one the directly-follows relations (edges) and their counts. Given a DFG, the probability of a specific trace is computed as the product of the transition (edge) probabilities along the trace. Each transition probability is calculated by dividing the count of the corresponding edge by the count of its source node. In the original ER implementation, stochastic process models are converted into stochastic deterministic finite automata (SDFAs) to compute trace probabilities. To replicate this in the DFG-based approach, special Beginning-Of-Sequence (BOS) and End-Of-Sequence (EOS) activities are introduced to all traces and included in the DFG. Incorporating these artificial transitions effectively replicates the initialization and termination probabilities that would otherwise be handled in an SDFA. The traces with the added BOS and EOS activities are also the ones used to discover the DFG, i.e., the DFGs will also include those nodes. Since all traces now start with BOS and end with EOS, the transition probabilities from the BOS node to any other node can be seen as initialization probabilities, and those from each node to EOS as termination probabilities.
@@
0: Variant log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$: $K$ distinct variants $t_{i}$, each with multiplicity $m_{i}$ directly-follows graph $G=(N,E,c_{N},c_{E})$: with $N=$ set of activity labels incl. BOS and EOSE=E= set of directed edges $c_{N}$ and $c_{E}$ the count of occurrences of the nodes and edges (the DFG). underflow threshold $\varepsilon$
@@
[…]
@@
[…]
@@
[…]
@@
Our (simplified) ER calculation is formalized in Algorithm 1. The procedure begins by calculating the total number of traces $T$ based on the multiplicities of the variants (Line 1) and initializing the total ER accumulator (Line 2). Next, the algorithm precomputes the transition probability for every edge in the DFG (Lines 3–5); this is done by dividing the count of a specific edge by the count of its source node. The algorithm then iterates through every distinct trace variant $t_{i}$ in the log (Line 6). As noted above, each trace is treated as a sequence starting with BOS and ending with EOS (Line 7). We initialize the probability $p$ for the current trace (Line 8) and calculate its total probability by multiplying the transition probabilities of every consecutive pair of activities along the path (Lines 9–11). To handle numerical instability, we safeguard the calculated probability against underflow by ensuring it is at least $\varepsilon$ (Line 12). We then calculate the information cost (ER) and add it to the running sum (Line 13).
Usually, ER of a log-model combination is calculated as the total ER (over all traces) [Alkhammash2022]. To be able to be able to compare ER scores across different clusters (or event logs) with different amounts of traces, we also opted to use the average ER score over traces (Line 15). In the clustering algorithms, we calculate the ER value of individual traces only, so there is no difference. However, in the full results (see A), we report both the average ER and the full sum.
Additionally, as mentioned above, the original ER formulation also handles non-fitting traces using a background model. However, in our method, ER scores are computed only for trace–DFG combinations in which the trace was used to discover the unfiltered DFG. The model therefore always fits the trace, and its probability can be directly derived from the DFG without resorting to a background model. Finally, as mentioned before, to avoid numerical underflow when computing probabilities for complex models (e.g., traces with many loops), our implementation enforces a lower bound on trace probabilities. Specifically, each probability is constrained to be at least $\varepsilon=10^{-10}$ before the logarithmic transformation. This effectively introduces a ceiling on the information cost for traces with extremely low likelihood. While this may yield slightly optimistic absolute ER scores for very complex (sub)processes, we argue that for the purpose of clustering, where the focus lies on the interpretability of the process models in each cluster, distinguishing between “extremely unlikely” and “infinitesimally unlikely” traces provides little additional value.
@@ 4.2 Entropic Clustering
Since ER computation scales linearly with input size [Alkhammash2022], the clustering process (excluding initialization) is also linear in complexity with the exact runtime depending on trace lengths, the preselected number of clusters, and the number of unique variants. An alternative approach, which adds the variant to the cluster for which the total cluster ER score is minimal (after adding), rather than the specific variant’s ER, was also explored. However, initial results showed inferior performance, leading us to omit it from further experiments.
@@
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$, where $t_{i}$ is a trace variant and $m_{i}$ its multiplicity number of clusters $k$ underflow threshold $\varepsilon$
@@
5:  Initialize DFG $G_{j}=(N,E,c_{N},c_{E})$ from seed $s_{j}$:
6:    $N\leftarrow$ Activity labels in $L\cup\{\mathrm{BOS},\mathrm{EOS}\}$
7:    $E\leftarrow$ Directed edges observed in $L$
8:    $c_{N},c_{E}\leftarrow$ Counts of nodes/edges in $s_{j}$ (starting counts)
9:end for
10:for each variant $v\in\mathcal{V}\setminus\{s_{1},\ldots,s_{k}\}$, in decreasing order of multiplicity do
11:  Let $L_{temp}\leftarrow\{(v,1)\}$ {Prepare single-instance log for ER calc}
12:  for each cluster $C_{j}\in\mathcal{C}$ do
13:   Temporarily update $G_{j}$ for variant $v$:
14:     Increment $c_{N}$ and $c_{E}$ by 1 for all nodes/edges in path of $v$
15:   $\text{score}_{j}\leftarrow\textsc{ComputeER}(L_{temp},G_{j},\varepsilon)$ {Call Algorithm 1}
16:   Revert changes to $G_{j}$ (subtract 1 from relevant counts)
17:  end for
18:  Let $j^{*}=\arg\min_{j}\text{score}_{j}$
19:  Assign $v$ to cluster $C_{j^{*}}$
20:  Update $G_{j^{*}}$ permanently (add actual multiplicity $m_{v}$ to counts)
21:end for
22:return Final cluster assignment $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$
Figure 1: Illustrative example of how the method selects the cluster.
In addition to the standard algorithm, we also investigate a recursive, hierarchical variant of Entropic Clustering, denoted as ECsplit. This approach is detailed in Algorithm 3. Instead of initializing $k$ clusters simultaneously, ECsplit begins with a single cluster containing the entire event log. It then proceeds iteratively: in each step, the algorithm identifies the “worst-performing” cluster—defined as the cluster with the highest total (or average) Entropic Relevance score. This high-entropy cluster is then split into two sub-clusters using the standard Entropic Clustering procedure with $k=2$. This process repeats until the pre-defined total number of clusters is reached. The theoretical motivation for this greedy strategy is drawn from recursive partitioning methods like ID3 for decision trees [Quinlan1986]. The hypothesis is that by focusing computational effort on refining the most unstructured subsets of traces (those with the poorest model fit), the algorithm might converge to a better global solution or achieve comparable results with lower computational overhead. We evaluate this hypothesis against the standard global initialization in our experimental section.
Algorithm 3 Entropic Clustering Split (ECsplit)
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$ desired number of clusters $k$ underflow threshold $\varepsilon$
@@
1: Initialize clustering $\mathcal{C}\leftarrow\textsc{EC}(L,2,\varepsilon)$ {Call Entropic Clustering (Alg. 2)}
@@
4:   Let $L_{j}\subseteq L$ be the sub-log containing variants assigned to $C_{j}$
5:   Build DFG $G_{j}$ from $L_{j}$ (calculating $N,E,c_{N},c_{E}$)
6:   $\text{score}_{j}\leftarrow\textsc{ComputeER}(L_{j},G_{j},\varepsilon)$ {Calculate ER (Alg. 1)}
7:  end for
8:  Let $j^{*}=\arg\max_{j}\text{score}_{j}$ {Identify cluster with highest entropy}
9:  Let $L_{j^{*}}$ be the variant log of cluster $C_{j^{*}}$
10:  $\{C_{new1},C_{new2}\}\leftarrow\textsc{EC}(L_{worst},2,\varepsilon)$ {Split the worst cluster}
11:  Update $\mathcal{C}\leftarrow(\mathcal{C}\setminus\{C_{j^{*}}\})\cup\{C_{new1},C_{new2}\}$
12:end while
13:return Final cluster assignment $\mathcal{C}$